# The Oracle Complexity of Simplex-based Matrix Games: Linear Separability and Nash Equilibria

### v1

@@ Abstract
We study the problem of solving matrix games of the form $\max_{\mathbf{w}\in\mathcal{W}}\min_{\mathbf{p}\in\Delta}\mathbf{p}^{\top}A\mathbf{w}$, where $A$ is some matrix and $\Delta$ is the probability simplex. This problem encapsulates canonical tasks such as finding a linear separator and computing Nash equilibria in zero-sum games. However, perhaps surprisingly, its inherent complexity (as formalized in the standard framework of oracle complexity (Nemirovski and Yudin, 1983)) is not well-understood. In this work, we first identify different oracle models which are implicitly used by prior algorithms, amounting to multiplying the matrix $A$ by a vector from either one or both sides. We then prove complexity lower bounds for algorithms under both access models, which in particular imply a separation between them. Specifically, we start by proving that algorithms for linear separability based on one-sided multiplications must require $\Omega(\gamma_{A}^{-2})$ iterations, where $\gamma_{A}$ is the margin, as matched by the Perceptron algorithm. We then prove that accelerated algorithms for this task, which utilize multiplications from both sides, must require $\tilde{\Omega}(\gamma_{A}^{-2/3})$ iterations, establishing the first oracle complexity barrier for such algorithms. Finally, by adapting our lower bound to $\ell_{1}$ geometry, we prove that computing an $\epsilon$-approximate Nash equilibrium requires $\tilde{\Omega}(\epsilon^{-2/5})$ iterations, which is an exponential improvement over the previously best-known lower bound due to Hadiji et al. (2024).
@@ 1 Introduction
Perhaps surprisingly, although these are canonical problems with a long history (see Sec. 3), the inherent complexity of solving them is relatively little studied. The goal of this paper is to study this question through the standard framework of oracle complexity (Nemirovski and Yudin, 1983). Specifically, we are interested in the performance limits of iterative algorithms, where each iteration is based on a simple computation involving the matrix $A$, such as multiplying it by some vector or extracting a row of $A$. This interaction between the algorithm and the matrix can be modeled as accessing an oracle, which simulates this computation and provides the algorithm with the result. We can then ask and rigorously study how many such oracle queries/computations are required, so that an algorithm with no prior knowledge of $A$ will solve the associated matrix problem up to a given level of precision (see Sec. 2 for more details). In this paper, we focus on the high-dimensional setting, where the size of the matrix $A$ is essentially unrestricted, and we are interested in bounds which are independent of (or only weakly dependent on) the matrix size.
@@
Considering linear separability, it is well known that for matrices with normalized rows, the Perceptron algorithm finds a linear separator using $O(\gamma_{A}^{-2})$ iterations (each involving a single matrix-vector multiplication and extraction of a single row of $A$), independently of $n,d$ (Novikoff, 1962). Half a century later, Soheili and Pena (2012); Yu et al. (2014) provided accelerated algorithms, which find a separator using only $O(\sqrt{\log n}\cdot\gamma_{A}^{-1})$ iterations. However, a closer inspection of these methods reveal that they rely on a stronger oracle access to the matrix $A$, amounting to multiplying $A$ by vectors on both sides (instead of just one-sided multiplications as required by the Perceptron algorithm). Thus, one may ask whether such two-sided operations are necessary for these accelerated results, and how close they are to being optimal.
@@
As for approximating Nash equilibria, the best current result we are aware of is due to Rakhlin and Sridharan (2013), who provided an algorithm which returns an $\epsilon$-suboptimal solution [^1] More precisely, the guarantee is that the algorithm returns an $\epsilon$-approximate saddle point, namely $(\hat{\mathbf{w}},\hat{\mathbf{p}})\in\Delta^{d-1}\times\Delta^{n-1}$ such that $\max_{\mathbf{w}\in\Delta^{d-1}}\hat{\mathbf{p}}^{\top}A\mathbf{w}-\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\hat{\mathbf{w}}\leq\epsilon$. This implies $\epsilon$-suboptimality, since $\max\{\max_{\mathbf{w}\in\Delta^{d-1}}\hat{\mathbf{p}}^{\top}A\mathbf{w}-\gamma_{A},~{}\gamma_{A}-\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\hat{\mathbf{w}}\}\leq\epsilon$ where $\gamma_{A}$ is the optimal value. Since the latter inequality also implies the former (with $2\epsilon$ replacing $\epsilon$), we see that the two notions are the same up to a factor of 2. using $O(\log(n)\cdot\epsilon^{-1})$ two-sided matrix multiplication queries. Corresponding lower bounds were missing, and only recently Hadiji et al. (2024) proved a lower bound of $\Omega(\log(1/n\epsilon))$ for sufficiently small $\epsilon=\mathrm{poly}(1/n)$ and $n=d$.
@@
  • One-sided vs. two-sided oracles (Sec. 3:) We start by identifying and formalizing different oracle models for matrix games that are implicitly used by existing algorithms, as we will see that these different oracles lead to different complexities. One oracle model corresponds to querying rows of $A$, together with “one-sided” multiplication queries of the form $\mathbf{w}\mapsto A\mathbf{w}$. The second (and stronger) oracle model we will consider allows “two-sided” multiplications $(\mathbf{p},\mathbf{w})\mapsto(\mathbf{p}^{\top}A,A\mathbf{w})$.
  • Linear separability with a one-sided oracle (Theorem 4.1): We first prove that any deterministic algorithm, that performs row queries and one-sided multiplication queries, must require $\Omega(\gamma_{A}^{-2})$ queries in the worst-case in order to find a linear separator. In particular, this establishes the optimality of the Perceptron algorithm under this oracle model.
  • Linear separability with a two-sided oracle (Theorem 4.2): We prove that any deterministic algorithm, which performs two-sided multiplication queries must require $\tilde{\Omega}(\gamma_{A}^{-2/3})$ queries in the worst-case in order to find a linear separator (where the $\tilde{\Omega}$ hides a logarithmic dependence on the matrix size). To the best of our knowledge, this is the first oracle complexity lower bound for linear separability, which applies even to accelerated algorithms. Compared to the lower bound for the one-sided oracle, the proof of the lower bound for the two-sided oracle is substantially more involved, and requires some new proof ideas.
@@
We conclude with a discussion and directions for future work in Sec. 6.
@@ 2 Preliminaries
In this paper, we focus on deterministic algorithms, whose oracle queries and output are deterministic function of the previous oracle responses. Since all state-of-the-art algorithms for the problems we consider are deterministic, this is without too much loss of generality. However, extending our lower bounds to randomized algorithms is certainly an interesting direction for future work (see Sec. 6 for more details).
@@ 3 Oracle models
Thus, the Perceptron’s convergence guarantee implies that $O(\gamma_{A}^{-2})$ queries to a one-sided oracle is sufficient for finding a linear separator.
In fact, a more general result can be achieved by applying the well-known subgradient method on the equivalent convex problem $\min_{\mathbf{w}:\|\mathbf{w}\|\leq 1}\max_{\mathbf{p}\in\Delta}\mathbf{p}^{\top}(-A)\mathbf{w}$. Specifically, by standard results, this method requires $O(1/\epsilon^{2})$ subgradient computations in order to find a vector $\mathbf{w}$ whose value is $\epsilon$-suboptimal (Nesterov et al., 2018). This holds for any $\epsilon$, whereas the guarantee for the Perceptron is merely for the special case $\epsilon=\gamma_{A}$ (namely, we seek a solution whose value is $>0$, with the optimal value being $\gamma_{A}$). From an oracle-complexity perspective, implementing such methods requires access to supergradients of the function $f(\mathbf{w})=\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}=\min_{l\in[n]}(A\mathbf{w})_{l}$, which equal $A_{l_{\min}}$ where $l_{\min}\in\arg\min_{l\in[n]}(A\mathbf{w})_{l}$. Thus, we can model these methods as iteratively interacting with the following supergradient oracle:
@@
Given some $\mathbf{w}\in\mathbb{R}^{d}$, the oracle $\mathcal{O}_{\partial}^{A}(\mathbf{w})$ returns $A_{l_{\min}}$ where $l_{\min}\in\arg\min_{l\in[n]}(A\mathbf{w})_{l}$
@@
Clearly, a two-sided oracle is stronger than a one-sided oracle, since $\mathcal{O}_{1}^{A}(l,\mathbf{w})$ can be simulated by $\mathcal{O}_{2}^{A}(\mathbf{e}_{l},\mathbf{w})$. As far as we know, all existing accelerated algorithms for linear separability can be implemented with such an oracle, so any lower bound for algorithms based on this oracle will apply to them.
Although we have considered so far the linear separability problem, a two-sided oracle is also very natural to model algorithms for other matrix games. In particular, for computing a Nash equilibrium, a two-sided oracle corresponds precisely to a first-order (or gradient) oracle, which given $\mathbf{w},\mathbf{p}$, returns the gradient of the function $(\mathbf{w},\mathbf{p})\mapsto\mathbf{p}^{\top}A\mathbf{w}$ (namely $\mathbf{p}^{\top}A$ and $A\mathbf{w}$). Under this well-studied setting, the best-known method due to Rakhlin and Sridharan (2013) requires $O(\log(n)/\epsilon)$ two-sided oracle calls in order to find an $\epsilon$-optimal solution (see Hadiji et al. (2024) for a detailed discussion of other results). However, very little work appears to exist on lower bounds, with the notable exception of Hadiji et al. (2024), whose bound is $\Omega(\log(1/(n\epsilon)))$ – namely, logarithmic in $1/\epsilon$ – and when $n=d$.
@@
In all oracle definitions, we do not restrict the input $\mathbf{w}$ to lie in $\mathcal{W}$, nor $\mathbf{p}$ to lie in $\Delta^{n-1}$. Thus, our lower bounds will apply equally to algorithms which query inside or outside these domains.
@@
Dimension-dependent oracle complexity for linear separability.
Besides the accelerated algorithms discussed earlier (with convergence rate $O(\log(n)\cdot\epsilon^{-1})$ for linear separability), there exists another family of rescaling-based methods, which can achieve an even faster (logarithmic) dependence on $\epsilon^{-1}$, but with iteration bounds scaling polynomially with the matrix dimensions (Dunagan and Vempala, 2004; Belloni et al., 2009; Pena and Soheili, 2016; Dadush et al., 2020). This is akin to the situation in convex optimization, where one can trade off between algorithms with dimension-independent, $\text{poly}(\epsilon^{-1})$-dependent guarantees (using gradient or subgradient methods), and algorithms with poly-dimension-dependent, $\log(\epsilon^{-1})$-dependent guarantees (using methods such as interior points, ellipsoids, or center-of-gravity, see (Nesterov et al., 2018)). Since our focus is on size-independent (or near independent) bounds, these family of methods are orthogonal to our work, although understanding the ultimate limits in that regime is interesting as well.
@@ 4.1 Oracle complexity with a one-sided Oracle
Proof of Thm. 4.1.
The proof is directly inspired by standard oracle complexity lower bounds for convex Lipschitz optimization (cf. Nesterov et al., 2018), as well as the standard lower bound proof of the Perceptron algorithm (cf. Shalev-Shwartz and Ben-David, 2014, section 9.5, question 3). The main idea is to construct $A$’s row as mutually orthogonal unit vectors, which are also orthogonal to the algorithm’s queries $\mathbf{w}_{t}$. Therefore, the algorithm can recover the rows one at a time, by querying for a particular row. However, if the number of rows is larger than $T$, there will remain rows orthogonal to $\mathbf{w}_{T+1}$. Therefore, $\mathbf{w}_{T+1}$ will not be a linear separator for $A$. On the other hand, since $A$ has $T+1$ orthogonal rows, it can be shown to be linearly separable with margin $1/\sqrt{T+1}$.
@@
On the other hand, the unit vector $\mathbf{w}:=\frac{1}{\sqrt{T+1}}\sum_{l=1}^{T+1}A_{l}$ (where $A_{l}$ is the $l$-th row of $A$, with the rows being mutually orthogonal unit vectors) satisfies
$A\mathbf{w}~{}=~{}\frac{1}{\sqrt{T+1}}\cdot\mathbf{1}~{}.$
Therefore, $\max_{\mathbf{w}:\|\mathbf{w}\|\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}\geq\frac{1}{\sqrt{T+1}}$. ∎

### v2

@@ Abstract
We study the problem of solving matrix games of the form $\max_{\mathbf{w}\in\mathcal{W}}\min_{\mathbf{p}\in\Delta}\mathbf{p}^{\top}A\mathbf{w}$, where $A$ is some matrix and $\Delta$ is the probability simplex. This problem encapsulates canonical tasks such as finding a linear separator and computing Nash equilibria in zero-sum games. However, perhaps surprisingly, its inherent complexity (as formalized in the standard framework of oracle complexity (Nemirovski and Yudin, 1983)) is not well-understood. In this work, we first identify different oracle models which are implicitly used by prior algorithms, amounting to multiplying the matrix $A$ by a vector from either one or both sides. We then prove complexity lower bounds for algorithms under both access models, which in particular imply a separation between them. Specifically, we start by showing that algorithms for linear separability based on one-sided multiplications must require $\Omega(\gamma_{A}^{-2})$ iterations, where $\gamma_{A}$ is the margin, as matched by the Perceptron algorithm. We then prove that accelerated algorithms for this task, which utilize multiplications from both sides, must require $\tilde{\Omega}(\gamma_{A}^{-2/3})$ iterations, establishing the first oracle complexity barrier for such algorithms. Finally, by adapting our lower bound to $\ell_{1}$ geometry, we prove that computing an $\epsilon$-approximate Nash equilibrium requires $\tilde{\Omega}(\epsilon^{-2/5})$ iterations, which is an exponential improvement over the previously best-known lower bound due to Hadiji et al. (2024).
@@ 1 Introduction
Perhaps surprisingly, although these are canonical problems with a long history (see Section 3), the inherent complexity of solving them is relatively little studied. The goal of this paper is to study this question through the standard framework of oracle complexity (Nemirovski and Yudin, 1983). Specifically, we are interested in the performance limits of iterative algorithms, where each iteration is based on a simple computation involving the matrix $A$, such as multiplying it by some vector or extracting a row of $A$. This interaction between the algorithm and the matrix can be modeled as accessing an oracle, which simulates this computation and provides the algorithm with the result. We can then ask and rigorously study how many such oracle queries/computations are required, so that an algorithm with no prior knowledge of $A$ will solve the associated matrix problem up to a given level of precision (see Section 2 for more details). In this paper, we focus on the high-dimensional setting, where the size of the matrix $A$ is essentially unrestricted, and we are interested in bounds which are independent of (or only weakly dependent on) the matrix size.
@@
Considering linear separability, it is well known that for matrices with normalized rows, the Perceptron algorithm finds a linear separator using $O(\gamma_{A}^{-2})$ iterations (each involving a single matrix-vector multiplication and extraction of a single row of $A$), independently of $n,d$ (Novikoff, 1962). Half a century later, Soheili and Pena (2012); Yu et al. (2014) used acceleration techniques due to Nesterov (2005a); Nemirovski (2004) respectively, and provided accelerated algorithms which find a separator using only $O(\sqrt{\log n}\cdot\gamma_{A}^{-1})$ iterations. However, a closer inspection of these methods reveal that they rely on a stronger oracle access to the matrix $A$, amounting to multiplying $A$ by vectors on both sides (instead of just one-sided multiplications as required by the Perceptron algorithm). Thus, one may ask whether such two-sided operations are necessary for these accelerated results, and how close they are to being optimal.
@@
As for approximating Nash equilibria, the best known algorithms return an $\epsilon$-suboptimal solution [^1] More precisely, these algorithms guarantee returning an $\epsilon$-approximate saddle point, namely $(\hat{\mathbf{w}},\hat{\mathbf{p}})\in\Delta^{d-1}\times\Delta^{n-1}$ such that $\max_{\mathbf{w}\in\Delta^{d-1}}\hat{\mathbf{p}}^{\top}A\mathbf{w}-\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\hat{\mathbf{w}}\leq\epsilon$. This implies $\epsilon$-suboptimality, since $\max\{\max_{\mathbf{w}\in\Delta^{d-1}}\hat{\mathbf{p}}^{\top}A\mathbf{w}-\gamma_{A},~{}\gamma_{A}-\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\hat{\mathbf{w}}\}\leq\epsilon$ where $\gamma_{A}$ is the optimal value. Since the latter inequality also implies the former (with $2\epsilon$ replacing $\epsilon$), we see that the two notions are the same up to a factor of 2. using $O(\sqrt{\log(n)\log(d)}\cdot\epsilon^{-1})$ two-sided matrix multiplication queries, a rate achieved by several different algorithms in the literature (Nemirovski, 2004; Nesterov, 2005b, 2007; Rakhlin and Sridharan, 2013). Corresponding lower bounds were missing, and only recently Hadiji et al. (2024) proved a lower bound of $\Omega(\log(1/n\epsilon))$ for sufficiently small $\epsilon=\mathrm{poly}(1/n)$ and $n=d$.
@@
  • One-sided vs. two-sided oracles (Section 3:) We start by identifying and formalizing different oracle models for matrix games that are implicitly used by existing algorithms, as we will see that these oracles lead to different complexities. One oracle model corresponds to querying rows of $A$, together with “one-sided” multiplication queries of the form $\mathbf{w}\mapsto A\mathbf{w}$. The second (and stronger) oracle model we will consider allows “two-sided” multiplications $(\mathbf{p},\mathbf{w})\mapsto(\mathbf{p}^{\top}A,A\mathbf{w})$.
  • Linear separability with a one-sided oracle (Theorem 4.1): We first show that any deterministic algorithm that performs row queries and one-sided multiplication queries, must require $\Omega(\gamma_{A}^{-2})$ queries in the worst-case in order to find a linear separator. The claim is proved using a classic lower bound technique due to Nemirovski and Yudin (1983). In particular, this establishes the optimality of the Perceptron algorithm under this oracle model.
  • Linear separability with a two-sided oracle (Theorem 4.2): We prove that any deterministic algorithm which performs two-sided multiplication queries, must require $\tilde{\Omega}(\gamma_{A}^{-2/3})$ queries in the worst-case in order to find a linear separator (where the $\tilde{\Omega}$ hides a logarithmic dependence on the matrix size). To the best of our knowledge, this is the first oracle complexity lower bound for linear separability, which applies even to accelerated algorithms. Compared to the lower bound for the one-sided oracle, the proof of the lower bound for the two-sided oracle is substantially more involved, and requires some new proof ideas.
@@
We conclude with a discussion and directions for future work in Section 6.
@@ 2 Preliminaries
In this paper, we focus on deterministic algorithms, whose oracle queries and output are deterministic function of the previous oracle responses. Since all state-of-the-art algorithms for the problems we consider are deterministic, this is without too much loss of generality. However, extending our lower bounds to randomized algorithms is certainly an interesting direction for future work (see Section 6 for more details).
@@ 3 Oracle models
Thus, the Perceptron’s convergence guarantee implies that $O(\gamma_{A}^{-2})$ queries to a one-sided oracle is sufficient for finding a linear separator. In fact, a more general result can be achieved by applying the well-known subgradient method on the equivalent convex problem $\min_{\mathbf{w}:\|\mathbf{w}\|\leq 1}\max_{\mathbf{p}\in\Delta}\mathbf{p}^{\top}(-A)\mathbf{w}$. Specifically, by standard results, this method requires $O(1/\epsilon^{2})$ subgradient computations in order to find a vector $\mathbf{w}$ whose value is $\epsilon$-suboptimal (Nesterov, 2018). This holds for any $\epsilon$, whereas the guarantee for the Perceptron is merely for the special case $\epsilon=\gamma_{A}$ (namely, we seek a solution whose value is $>0$, with the optimal value being $\gamma_{A}$). From an oracle-complexity perspective, implementing such methods requires access to supergradients of the function $f(\mathbf{w})=\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}=\min_{l\in[n]}(A\mathbf{w})_{l}$, which equal $A_{l_{\min}}$ where $l_{\min}\in\arg\min_{l\in[n]}(A\mathbf{w})_{l}$. Thus, we can model these methods as iteratively interacting with the following supergradient oracle:
@@
Given some $\mathbf{w}\in\mathbb{R}^{d}$, the oracle $\mathcal{O}_{\partial}^{A}(\mathbf{w})$ returns $A_{l_{\min}}$ where $l_{\min}\in\arg\min_{l\in[n]}(A\mathbf{w})_{l}$.
@@
Clearly, a two-sided oracle is stronger than a one-sided oracle, since $\mathcal{O}_{1}^{A}(l,\mathbf{w})$ can be simulated by $\mathcal{O}_{2}^{A}(\mathbf{e}_{l},\mathbf{w})$. As far as we know, all existing accelerated algorithms for linear separability can be implemented with such an oracle, so any lower bound for algorithms based on this oracle will apply to them. We also note that lower bounds with respect to two-sided oracles were studied in the context of solving linear equations (Nemirovsky, 1992), which does not directly relate to our results or proofs.
Although we have considered so far the linear separability problem, a two-sided oracle is also very natural to model algorithms for other matrix games. In particular, for computing a Nash equilibrium, a two-sided oracle corresponds precisely to a first-order (or gradient) oracle, which given $\mathbf{w},\mathbf{p}$, returns the gradient of the function $(\mathbf{w},\mathbf{p})\mapsto\mathbf{p}^{\top}A\mathbf{w}$ (namely $\mathbf{p}^{\top}A$ and $A\mathbf{w}$). Under this well-studied setting, the best-known method dating back to Nemirovski (2004) requires $O(\sqrt{\log(n)\log(d)}/\epsilon)$ two-sided oracle calls in order to find an $\epsilon$-optimal solution (see also Nesterov (2005b, 2007); Rakhlin and Sridharan (2013) and Hadiji et al. (2024) for a detailed discussion of other results). However, very little work appears to exist on lower bounds, with the notable exception of Hadiji et al. (2024), whose bound is $\Omega(\log(1/(n\epsilon)))$ – namely, logarithmic in $1/\epsilon$ – and when $n=d$.
@@
In all oracle definitions, we do not restrict the input $\mathbf{w}$ to lie in $\mathcal{W}$, nor $\mathbf{p}$ to lie in $\Delta^{n-1}$. Thus, our lower bounds will apply equally to algorithms which can query outside these domains.
@@
Dimension-dependent oracle complexity.
Besides the accelerated algorithms for linear separability discussed earlier (with convergence rate $O(\log(n)\cdot\epsilon^{-1})$), there exists another family of rescaling-based methods, which can achieve an even faster (logarithmic) dependence on $\epsilon^{-1}$, but with iteration bounds scaling polynomially with the matrix dimensions (Dunagan and Vempala, 2004; Belloni et al., 2009; Pena and Soheili, 2016; Dadush et al., 2020). This is akin to the situation in convex optimization, where one can trade off between algorithms with dimension-independent, $\text{poly}(\epsilon^{-1})$-dependent guarantees (using gradient or subgradient methods), and algorithms with poly-dimension-dependent, $\log(\epsilon^{-1})$-dependent guarantees (using methods such as interior points, ellipsoids, or center-of-gravity, see Nesterov, 2018). Similarly, several works developed algorithms for general matrix games that reduce the $\epsilon$-dependence at the cost of polynomial dependencies on $n,d$, thus improving size-independent algorithms in certain parameter regimes (Carmon et al., 2019, 2024). Since our focus is on size-independent (or near independent) bounds, these family of methods are orthogonal to our work, although understanding the ultimate limits in those regimes is interesting as well.
@@ 4.1 Oracle complexity with a one-sided oracle
Proof of Theorem 4.1.
The proof is directly inspired by standard oracle complexity lower bounds for convex Lipschitz optimization due to Nemirovski and Yudin (1983), as well as the standard lower bound proof of the Perceptron algorithm (cf. Shalev-Shwartz and Ben-David, 2014, section 9.5, question 3). The main idea is to construct $A$’s row as mutually orthogonal unit vectors, which are also orthogonal to the algorithm’s queries $\mathbf{w}_{t}$. Therefore, the algorithm can recover the rows one at a time, by querying for a particular row. However, if the number of rows is larger than $T$, there will remain rows orthogonal to $\mathbf{w}_{T+1}$. Therefore, $\mathbf{w}_{T+1}$ will not be a linear separator for $A$. On the other hand, since $A$ has $T+1$ orthogonal rows, it can be shown to be linearly separable with margin $1/\sqrt{T+1}$.
@@
On the other hand, the unit vector $\mathbf{w}:=\frac{1}{\sqrt{T+1}}\sum_{l=1}^{T+1}A_{l}$ (where $A_{l}$ is the $l$-th row of $A$, with the rows being mutually orthogonal unit vectors) satisfies $A\mathbf{w}~{}=~{}\frac{1}{\sqrt{T+1}}\cdot\mathbf{1}$, and therefore,
$\max_{\mathbf{w}:\|\mathbf{w}\|\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~{}\geq~{}\frac{1}{\sqrt{T+1}}~{}.$